OpenStudy (anonymous):

Solve 3 (√x-3) – 5 = 16 The √x-3 isnt actually in parentheses i just didnt know how to make it where it didnt seem as the -5 was in there too. Sorry for any confusion.

OpenStudy (anonymous):

Choices: x = 5 x = 19 x = 52 x = 67

OpenStudy (anonymous):

i think its \[\sqrt{x-3}\]

OpenStudy (anonymous):

using ^, we have 3(sqrt of x-3) = 16+5 3. (sqrt of x-3) = 21 or (sqrt of x-3) = 7 Squaring both sides, x-3 = 49 x = 52.

OpenStudy (anonymous):

yea but i cant do it like that in the question

OpenStudy (anonymous):

\[\Large 3{\sqrt{x-3}}~~–~5 = 16\]

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

@touseii45 already solved look up

OpenStudy (anonymous):

oh lol i didnt even see that haha

OpenStudy (anonymous):

thanks(: